OpenStudy (anonymous):

The minute hand of the clock moves 1/10 degree every second. If you look @ the clock when the minute hand is 10 degrees past the 12, you can use the equation y= 1/16x+10 to find how many degrees past the 12 the minute hand is after x second. Graph the equation and tell whether it is a linear.

OpenStudy (anonymous):

it's linear based on the fact that y is based on y=mx+b which is linear

OpenStudy (anonymous):



OpenStudy (dumbcow):

shouldn't the equation be: y = 1/10x +10 since the rate the minute hand is moving is 1/10 degree per second

OpenStudy (anonymous):

@dumbcow I think you're on the right track, but how do I graph it ?

OpenStudy (dumbcow):

its in the form y=mx+b , where m is slope (rise/run) and b is y_intercept start at y_intercept, then plot points along the slope 1/10....up 1, over 10